MT&T Telecommunications Company charges $90 per month for unlimited use of cell phone to any number (cell or land-line). Thei steps and correct multiple choice answer.
Answer #1

RR = retention rate = 1- attrition rate

DR = Discount rate

CLV = Margin. RR/(1+DR-RR)

Old RR = 1- 1.8/100 = 1- 0.018 = 0.982

DR = 0.01

Revenue=$ 90

Costs/month:

1. Variable cost =$ 30

2. Marketing cost = $30 / 12 = $2.5

Old Margin(monthly) = Revenue(90) - variable cost(30) - marketing cost per month(2.5) = 90 - 30 - 2.5 = 57.5

Old CLV = 57.5 * 0.982 / (1 + 0.01 - 0.982) = 2016.61

New RR = 1-1.5/100 = 1- 0.015 = 0.985

DR = 0.01

Revenue=$ 90

Costs/month:

1. Variable cost =$ 30

2. Marketing cost = $30 / 12

3. Retention cost = $36 / 12 = $3

New Margin(monthly) = Revenue(90) - variable cost(30) - marketing cost per month(2.5) - retention cost(3)= 90 - 30 - 2.5 -3= 54.5

New CLV = 54.5 * 0.985 / (1 + 0.01 - 0.985) = 2147.3

retention spending increases by $3 + (0.985-0.982)*100 = $3

So, they shouldn't do it.